# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

### v5

@@ 1 Introduction
Magnetars are highly magnetized isolated neutron stars (Duncan & Thompson, 1992; van Kerkwijk et al., 1995; Kouveliotou et al., 1998; Banas et al., 1997; Kaspi & Beloborodov, 2017), and are characterized by their slow rotation period (P s), rapidly spinning down ( ), and relatively young age (typically about 1000 yr). We currently know about 30 magnetars (Olausen & Kaspi, 2014)111http://www.physics.mcgill.ca/~pulsar/magnetar/main.html. Magnetars are manifested themselves by persistent emission and emit bursts/flares simultaneously in the X-/gamma-ray band (Esposito et al., 2021).
@@
Based on their luminosity and duration, magnetar bursts can be divided into three classes (Woods & Thompson, 2006): Short-duration burst, which consists of single or multiple peaks, is the most typical magnetar burst with a burst duration 0.01 and the luminosity around ; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ( ) and peak luminosity around ; Giant flare, the rarest and the most powerful energetic burst is characterized by a significantly higher luminosity than a typical magnetar burst and a light curve with a hard initial spike and rapidly decaying tail modulated at the spin period of the magnetar.
@@
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by the Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos et al., 2014), while it is worth noting that Xie et al. (2022) found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan et al., 2009) before this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, and , respectively. This indicates the estimated magnetic field at the pole is (Israel et al., 2016). It has experienced multiple outbursts from 2014 to 2022 (Younes et al., 2017; Lin et al., 2020b, a; Cai et al., 2022a, b; Xie et al., 2022; Rehan & Ibrahim, 2023, 2024). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which burst forests were observed (Kaneko et al., 2021; Younes et al., 2020). Moreover, the source entered a new outburst (Borghese et al., 2022). Among these bursts there is an X-ray burst (Li et al., 2021; Mereghetti et al., 2020; Ridnaia et al., 2021; Tavani et al., 2020) associated with a fast radio burst, FRB 200428 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020a). Additional coincident radio and X-ray bursts were observed during the active episode of 2022 (Li et al., 2021; Xiao et al., 2022a; Zhang et al., 2023; Wang et al., 2024, 2022a; Li et al., 2022).
@@
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observations data from January 2021 to December 2022. As in Paper I of this series work for GECAM observations of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (in prep.), we will give a detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section 2, we present the observations, the search process, and report the burst catalog. In Section 3, we present the catalog analysis method and the results. Finally, a summary and discussions are given in Section 4.
@@
2 Observation and Burst Search
@@
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, Xiao et al., 2022b), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang et al., 2023), and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang et al., 2024). These four satellites are all-sky monitors with a large field of view that could monitor various gamma-ray transients except for the area blocked by Earth. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 charged particle detectors (CPDs), while GECAM-C has 12 GRDs and 2 CPDs. All GECAM-B’s GRDs and most GECAM-C’s GRDs operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The event-by-event (EVT) data from GECAM’s GRD detectors is used for further analysis. The GECAM achieves a high time resolution of 0.1 , with a dead time of 4 for normal events and exceeding 69 for overflow events (Liu et al., 2021). We note that in this work only GECAM-B/C data are used because of power supply issues that GECAM-A has not been able to observe the source yet (Li et al., 2022).
2.1 Search and Identification
@@
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan & Ibrahim, 2023, 2024; Wood & Fermi-GBM Team, 2022). A search pipeline developed by Cai et al. (2025) for gamma-ray bursts (GRBs) is used to do a blind search on GECAM daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B (2021-01 to 2022-01, Xie et al., 2022)222The 94 bursts shown in Xie et al. (2022) are the preliminary blind searched results that are reported in the Gamma-ray Coordinates Network (GCN) by the burst advocate (BA). We identify 82 bursts from SGR J1935+2154 using the targeted search in this paper..
@@
The search pipeline is implemented using a ground sensitive coherent search method that features a log-likelihood ratio (hereafter LR) comparing expected model counts to observed counts in each channel of each detector (Blackburn et al., 2015; Cai et al., 2025),
@@
where is the number of data sets in each channel and detector, is the total number of detectors and channels, is the background-subtracted counts and is the error of the expected counts (background+source), is the estimated background and is the error of the background counts, and represent the instrument response and the expected source amplitude (could estimated by a spectral model template), respectively. Therefore, the LR could be calculated with the response factors over the all-sky for a blind search or a certain area for a targeted search. Then, the continuous time-bins with the higher LR (higher than the LR threshold) could be considered as one signal (trigger/burst). The maximum LR gives the most possible spatial location of the source signal. The pipeline will mark the first time-bin, of which LR is higher than the LR threshold, as a trigger time. In this paper, the LR threshold is set to 20 and the search timescale (time-bin) is set to 10 ms.
@@
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold (typically as implemented in flight software), we conduct a targeted sub-threshold search for the area around the position of SGR J1935+2154 (R.A. and Decl. ) on GECAM daily EVT data333https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai et al. (2025). Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral model templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin et al., 2020b, a; Rehan & Ibrahim, 2023, 2024), we utilize the following spectral model templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameter setting is listed in Table 1 which is similar to Table 1 in Xie et al. (2022). The details of each template are presented below.
@@
Thousands of triggers (candidate bursts) have been found through the targeted search process. We perform a refined analysis for each candidate burst by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools444https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from the candidates should adhere to the following criteria (same as Xie et al., 2022):
@@
The burst location is within the error circle (at 3 ) of the source position (SGR J1935+2154, R.A. = 293.73°, Decl. = 21.90°, Israel et al., 2016);
@@
| All bursts | SoftBand | 34 | 24.19% |
| CPL | 74 | 60.60% | |
| OTTB | 79 | 59.33% | |
| Blackbody | 48 | 37.81% | |
| Powerlaw | 21 | 18.03% |
2.2 Search Results
A total of 159 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a bright radio burst FRB 20221014 (Dong & Chime/Frb Collaboration, 2022; Wang et al., 2022b; Frederiks et al., 2022; Maan et al., 2022; Giri et al., 2023)666We marker this radio burst as FRB 20221014 in this paper since the luminosity is several orders higher than the pulsed emission from SGR J1935+2154 (Zhu et al., 2023) although still a few orders lower than that of general FRBs.. It is worth noting that about one-third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM777Using the Fermi/GBM position history file (poshist) and the GBM data tool (Goldstein et al., 2022) to determine if SGR J1935+2154 is visible to Fermi/GBM or occulted by the Earth at these burst times..
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to the targeted search. The CPL, OTTB, and BB account for a proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai et al., 2021). As shown in Fig. 3, using a Powerlaw template to search for a magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing a much difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel (40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to found using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
@@ 3.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data of GECAM-B (before October 2022) and GECAM-C (after October 2022) to search periodicity by the Lomb-Scargle Periodogram method (Lomb, 1976; Scargle, 1982; VanderPlas, 2018). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (false alarm probability: 0.001, significance level ). The peaks around the 50 day are likely caused by observation windows or bursts gap (Xie et al., 2022).
@@
To uncover the burst activity of SGR J1935+2154 within a cycle, the start time of all bursts is folded into different phases at a given period day through the period folding method (CHIME/FRB Collaboration et al., 2020b; Zhang et al., 2021; Xie et al., 2022),
@@
where is the folded phase, is the burst start time, is the phase start point (MJD: 59215; UTC: 2021-01-01), and the floor is a function that returns the floor of the input number. As shown in Fig 7, the active day within a cycle (duty cycle) of SGR J1935+2154 bursts is 80%. The error of the period could be estimated to be 20 days with (Chime/Frb Collaboration et al., 2020),
@@
where is the active days in one cycle and is the longest time separation between burst arrival times. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is days, consistent with previous work (Xie et al., 2022, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
3.3 Burst Duration
@@
As previous studies (Lin et al., 2013), the magnetar burst duration is calculated with the Bayesian blocks algorithm (Scargle et al., 2013). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s (at least twice the spin period of SGR J1935+2154) are treated as background. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance larger than is considered to be part of the burst region. The first block with a significance higher than is marked as the burst start time (see Tables 3 and 4). As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple peaks (e.g., burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the duration of the entire burst is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
@@
The duration () of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We perform the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The mean values of burst duration for the four active episodes are ms for the first episode, ms for the second episode, ms for the third episode, and / ms (given by GECAM-B/C data) for the fourth active episodes. The full sample gives the mean value ms. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin et al., 2020b, a; Cai et al., 2022a; Rehan & Ibrahim, 2023, 2024).
@@
3.4 Burst Waiting Time

### v6

@@ I Introduction
Magnetars are highly magnetized isolated neutron stars (ThompsonDuncan1992; vanKerkwijk1995; Kouveliotou1998; Banas1997; KaspiBeloborodov2017), and are characterized by their slow rotation period (P s), rapidly spinning down ( ), and relatively young age (typically about 1000 yr). We currently know about 30 magnetars (Olausen2014)111http://www.physics.mcgill.ca/~pulsar/magnetar/main.html. Magnetars are manifested themselves by persistent emission and emit bursts/flares simultaneously in the X-/gamma-ray band (Esposito2021ASSL).
@@
Based on their luminosity and duration, magnetar bursts can be divided into three classes (WoodsThompson2006): Short-duration burst, which consists of single or multiple peaks, is the most typical magnetar burst with a burst duration 0.01 and the luminosity around ; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ( ) and peak luminosity around ; Giant flare, the rarest and the most powerful energetic burst is characterized by a significantly higher luminosity than a typical magnetar burst and a light curve with a hard initial spike and rapidly decaying tail modulated at the spin period of the magnetar.
@@
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by the Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos2014), while it is worth noting that Xie2022mnras found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan2009) before this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, and , respectively. This indicates the estimated magnetic field at the pole is (Israel2016). It has experienced multiple outbursts from 2014 to 2022 (Younes2017; Lin2020a; Lin2020b; Cai2022apjs_A; Cai2022apjs_b; Xie2022mnras; Rehan2023ApJ; Rehan2024ApJ). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which burst forests were observed (Kaneko2021ApJ; Younes2020ApJ). Moreover, the source entered a new outburst (Borghese2022MNRAS). Among these bursts there is an X-ray burst (Li2021Nat; Mereghetti2020; Ridnaia2021NatAs; Tavani2020ATel) associated with a fast radio burst, FRB 200428 (Bochenek2020; CHIMEFRBCollaboration2020a). Additional coincident radio and X-ray bursts were observed during the active episode of 2022 (LiWen2021; XiaoXiong2022mnras; Zhang2023nimpra; Wang2024ea; Wang2022ATel; Li2022ATel).
@@
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observations data from January 2021 to December 2022. As in Paper I of this series work for GECAM observations of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (in prep.), we will give a detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section II, we present the observations, the search process, and report the burst catalog. In Section III, we present the catalog analysis method and the results. Finally, a summary and discussions are given in Section IV.
@@
II Observation and Burst Search
@@
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, xiaoLiu2022mnras), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang2023nimpra), and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang2024ea). These four satellites are all-sky monitors with a large field of view that could monitor various gamma-ray transients except for the area blocked by Earth. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 charged particle detectors (CPDs), while GECAM-C has 12 GRDs and 2 CPDs. All GECAM-B’s GRDs and most GECAM-C’s GRDs operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The event-by-event (EVT) data from GECAM’s GRD detectors is used for further analysis. The GECAM achieves a high time resolution of 0.1 , with a dead time of 4 for normal events and exceeding 69 for overflow events (Liu2021arXiv). We note that in this work only GECAM-B/C data are used because of power supply issues that GECAM-A has not been able to observe the source yet (Li2021hzr).
II.1 Search and Identification
@@
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan2023ApJ; Rehan2024ApJ; Wood2022GCN). A search pipeline developed by Cai2024SCP for gamma-ray bursts (GRBs) is used to do a blind search on GECAM daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B (2021-01 to 2022-01, Xie2022mnras)222The 94 bursts shown in Xie2022mnras are the preliminary blind searched results that are reported in the Gamma-ray Coordinates Network (GCN) by the burst advocate (BA). We identify 82 bursts from SGR J1935+2154 using the targeted search in this paper..
@@
The search pipeline is implemented using a ground sensitive coherent search method that features a log-likelihood ratio (hereafter LR) comparing expected model counts to observed counts in each channel of each detector (Blackburn2015ApJS; Cai2024SCP),
@@
where is the number of data sets in each channel and detector, is the total number of detectors and channels, is the background-subtracted counts and is the error of the expected counts (background+source), is the estimated background and is the error of the background counts, and represent the instrument response and the expected source amplitude (could be estimated by a spectral model template), respectively. Therefore, the LR could be calculated with the response factors over the all-sky for a blind search or a certain area for a targeted search. Then, the continuous time-bins with the higher LR (higher than the LR threshold) could be considered as one signal (trigger/burst). The maximum LR gives the most possible spatial location of the source signal. The pipeline will mark the first time-bin, of which LR is higher than the LR threshold, as a trigger time. In this paper, the LR threshold is set to 20 and the search timescale (time-bin) is set to 10 ms.
@@
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold (typically as implemented in flight software), we conduct a targeted sub-threshold search for the area around the position of SGR J1935+2154 (R.A. and Decl. ) on GECAM daily EVT data333https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai2024SCP. Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral model templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin2020a; Lin2020b; Rehan2023ApJ; Rehan2024ApJ), we utilize the following spectral model templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameter setting is listed in Table 1 which is similar to Table 1 in Xie2022mnras. The details of each template are presented below.
@@
Thousands of triggers (candidate bursts) have been found through the targeted search process. We perform a refined analysis for each candidate burst by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools444https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from the candidates should adhere to the following criteria (same as Xie2022mnras):
@@
The burst location is within the error circle (at 3 ) of the source position (SGR J1935+2154, R.A. = 293.73°, Decl. = 21.90°, Israel2016);
@@
The solar flares, the particle events measured by CPD observations, and the Terrestrial gamma-ray flash (TGF, Briggs2013) and Terrestrial electron beam (TEB, Xiong2012) featured by being harder ( MeV) and shorter ( 1 ms) than magnetar bursts are excluded;
@@
| All bursts | SoftBand | 34 | 13.28% |
| CPL | 74 | 28.91% | |
| OTTB | 79 | 30.86% | |
| Blackbody | 48 | 18.75% | |
| Powerlaw | 21 | 8.20% |
II.2 Search Results
A total of 159 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a bright radio burst FRB 20221014 (DongChime2022ATel; WangXiong2022ATel; Frederiks2022ATel; Maan2022ATel; Giri2023arXiv)666We mark this radio burst as FRB 20221014 in this paper since the luminosity is several orders higher than the pulsed emission from SGR J1935+2154 (Zhu2023SciA) although still a few orders lower than that of general FRBs.. It is worth noting that about one-third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM777Using the Fermi/GBM position history file (poshist) and the GBM data tool (GbmDataTools) to determine if SGR J1935+2154 is visible to Fermi/GBM or occulted by the Earth at these burst times..
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to the targeted search. The CPL, OTTB, and BB account for a higher proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai2021mnras). As shown in Fig. 3, using a Powerlaw template to search for a magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing a large difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel (40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to find using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
@@ III.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data of GECAM-B (before October 2022) and GECAM-C (after October 2022) to search periodicity by the Lomb-Scargle Periodogram method (Lomb1976apss; Scargle1982apj; VanderPlas2018apjs). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (false alarm probability: 0.001, significance level ). The peaks around the 50 day are likely caused by observation windows or bursts gap (Xie2022mnras).
@@
To uncover the burst activity of SGR J1935+2154 within a cycle, the start time of all bursts is folded into different phases at a given period day through the period folding method (CHIMEFRB2020a; Zhang2021apj; Xie2022mnras),
@@
where is the folded phase, is the burst start time, is the phase start point (MJD: 59215; UTC: 2021-01-01), and the floor is a function that returns the floor of the input number. As shown in Fig 7, the active day within a cycle (duty cycle) of SGR J1935+2154 bursts is 80%. The error of the period could be estimated to be 20 days with (ChimeFrb2020Natur),
@@
where is the active days in one cycle and is the longest time separation between burst arrival times. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is days, consistent with previous work (Xie2022mnras, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
III.3 Burst Duration
@@
As previous studies (Lin2013ApJ), the magnetar burst duration is calculated with the Bayesian blocks algorithm (Scargle2013ApJ). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s (at least twice the spin period of SGR J1935+2154) are treated as background. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance larger than is considered to be part of the burst region. The first block with a significance higher than is marked as the burst start time (see Tables 3 and 4). As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple peaks (e.g., burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the duration of the entire burst is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
@@
The duration () of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We perform the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The mean values of burst duration for the four active episodes are ms for the first episode, ms for the second episode, ms for the third episode, and / ms (given by GECAM-B/C data) for the fourth active episodes. The full sample gives the mean value ms. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin2020a; Lin2020b; Cai2022apjs_A; Rehan2023ApJ; Rehan2024ApJ).
@@
III.4 Burst Waiting Time